OpenStudy (anonymous):

Length of a retangel is 4 times the width. the are of retangele is 64 square meters. Calc length and width of retangle.

OpenStudy (anonymous):

length=16m width=4m

OpenStudy (saifoo.khan):

Do you get how to do it?

OpenStudy (anonymous):

Hi but how did you get that?

OpenStudy (saifoo.khan):

Hello. I dont get it, he gets it. :p

OpenStudy (anonymous):

Arnab09 can you please assist with explanation

OpenStudy (saifoo.khan):

Area = L times W. Wdith = x Length = 4x. Area = L times W. 64 = 4x (x) 64 = 4x^2 x^2 = 16 x = 4, x = -4.

OpenStudy (saifoo.khan):

So, width = 4. Length = 4(x) = 4(4) = 16

OpenStudy (anonymous):

saifoo, dont joke... u cant get how i got that answer?? :(

OpenStudy (anonymous):

Thanks but still a little confused

OpenStudy (anonymous):

why??

OpenStudy (saifoo.khan):

lol, ok.. i wont joke.. ;)

OpenStudy (anonymous):

No i needed more time to unravel the message, but at now I've got it. thanks genius

OpenStudy (saifoo.khan):

NP.

OpenStudy (anonymous):

saifoo could you please assist with the ollowing pV(n-1) = c --in brackets on top next to v 2.8 ( 2x+1) x 9.4 (2x-1) = 6.3 (4x) - in brackets on tope next to last numbers

OpenStudy (saifoo.khan):

Sorry, can you elaborate the questions more?

OpenStudy (anonymous):

last question solvve x

OpenStudy (saifoo.khan):

\[2.8 ( 2x+1) \times 9.4 (2x-1) = 6.3 (4x)\]Something like that?

OpenStudy (anonymous):

Yes, only the brackets on top next to 8, 4 and 3

OpenStudy (anonymous):

first q just states pV (9n-1) = c on top next to V without brackets is n-1 this apply to first question as well

OpenStudy (saifoo.khan):

Umm, sorry.. i can't help you with those. :(

OpenStudy (anonymous):

thanks for the effort, will find a way to properly record this in order tomake it more clear. thanks anyhow

OpenStudy (saifoo.khan):

you welcome. :]